3. A psychologist believes that it will take at least an hour for certain disturbed


children to learn a task. A random sample of 30 of these children results in a


mean of 50 minutes to learn the task. Should the psychologist modify her belief


at the 0.01 level if the population standard deviation can be assumed to be 15


minutes?


1
Expert's answer
2022-05-12T03:23:42-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge 60"

"H_a:\mu<60"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a left-tailed test is "z_c = -2.3263."

The rejection region for this left-tailed test is "R = \{z: z < -2.3263\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{50-60}{15/\sqrt{30}}\approx-3.6515"

Since it is observed that "z = -3.6515 <-2.3263= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(Z<-3.6515)=0.00013," and since "p=0.00013<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than 60, at the "\alpha = 0.01" significance level.